OpenStudy (anonymous):

write 10m^15/2 in radical form

OpenStudy (anonymous):

\[10\sqrt{m^{15}}\]

OpenStudy (anonymous):

unless it was \[(10m)^{\frac{15}{2}}\]

OpenStudy (anonymous):

now that is radical form, not "simplest radical form" in which case the answer would be \[10m^7\sqrt{m}\]